You own an oil well that can produce 100k barrels of oil one year from now. Next year, the price of oil will be either $80/barrel if the economy is good and $40/barrel if the economy is bad. Oil costs $50/barrel to extract. One year forwards for oil has a forward price of $55 and the risk-free rate is 4%. What is the value of the oil well if you have to decide today whether to drill for oil or not? b.) Now suppose you can wait until the out come of the economy is decided before deciding to drill for oil or not. What is the new value of the oil well?
Answer #1

a) As the probabilities of good and bad economy is not given expected price of Oil can only be determined by the forward price

If the forward price is locked, profit after one year = $55 -$50 = $ 5 per barrel

So Value of Oil well today = Present value of profit after one year / barrel * no. of barrels

= 5 * e-0.04 * 100,000 barrels

= 4.8039472 * 100000

= $480,394.72

b) If we can wait until the outcome is assured then, we can start operations only when economy is good and we will get profit of $80 -$50 = $30 per barrel

So Value of Oil well today = Present value of profit after one year / barrel * no. of barrels

= 30 * e-0.04 * 100,000 barrels

= 28.8236832 * 100000

= $ 2,882,368.32